A thermometer reading of 100 degrees C was plunged into a tub of frozen water. If the thermometer reads 70C after 5 seconds, what will be the reading after 15 seconds?



1
Expert's answer
2021-12-21T16:51:04-0500

Let "T(t)" be the thermometer reading after "t" seconds. The surrounding temperature is "T_s=0\degree C," so Newton’s Law of Cooling states that


"\dfrac{dT}{dt}=k(T-T_s)"

"\dfrac{dT}{dt}=kT"

"\dfrac{dT}{T}=kdt"

"T(t)=Ce^{kt}"

"T(0)=C=100"

"T(t)=100e^{kt}"

"T(5)=100e^{k(5)}=70"

"e^{5k}=0.7"

"T(15)=100e^{k(15)}=100(e^{5k})^3"

"=100(0.7)^3=34.3(\degree C)"

The reading will be "34.3\degree C" after 15 seconds.